OpenStudy (anonymous):

Determine the slope between the points (-3, 0) and (0, 5) A. 3/5 B. 5/3 C. -5/3 D. -3/5 Determine the slope between the points (-1, -5) and (5, 4) A. 2/3 B. -1/4 C. 1/4 D. 3/2 Determine the slope between the points (-1, 0) and (-2, 3) A. 3 B. -1/3 C. -3 D. 1/3

OpenStudy (jdoe0001):

\(\large slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

all three of these problems can be solved by the slope formula. I will do the first one, but the rest is up to you... slope(m) = (y2 - y1) / (x2 - x1) (-3,0) x1 = -3 and y1 = 0 (0,5) x2 = 0 and y2 = 5 now lets sub our info into the formula... m = (5 - 0) / (0 - (-3) m = 5/(0 + 3) m = 5/3 It is that simple.

OpenStudy (anonymous):

okay thanks. lol. i understand it now

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

would the third one be 1/3 or just 3?

OpenStudy (anonymous):

m = (3 - 0) / (-2 - (-1) m = 3/(-2 + 1) m = 3/-1 m = -3